OpenStudy (anonymous):

x^2+6x+11 can be as (x+p)^2+q where p and q are integers to be found

Directrix (directrix):

The task is to create a perfect square trinomial using the given terms: x^2+6x+11 = (x^2+6x + 9) + 11 - 9 = ( x + 3) ^ 2 + 2 so p = 3 and q = 2 To get the 9, take half of 6 and square it. To maintain equality, subtract it after forming the perfect square trinomial.

OpenStudy (anonymous):

thank you my dear

Directrix (directrix):

Glad to help

OpenStudy (anonymous):

y-3x+2=0 y^2-x-6x^2=0 a simultanious equation